# IMAS2: Joint Agent Selection and Information-Theoretic Coordinated Perception In Dec-POMDPs

### v1

@@ Abstract.
Decentralized Partially Observable Markov Decision Process (Dec-POMDP), Agent Selection, Active Perception, Submodular Optimization.
@@ 1. Introduction
Autonomous multi-agent systems are increasingly deployed in environments where agents must actively gather information under uncertainty. Applications include teams of robots for surveillance and search-and-rescue Rybski et al. (2000); Kantor et al. (2006); Rosenfeld et al. (2015), sensor networks for target tracking He et al. (2006); Li et al. (2020), and cooperative perception in autonomous driving and aerial monitoring Xu et al. (2020); Wang et al. (2023). In such settings, an active perception agent must decide not only how to act, but also what to sense, given the past observation. In multi-agent cooperative active perception, it must additionally account for its knowledge of other agents’ possible observations to avoid redundancy and achieve better coordination. Meanwhile, in addition to perception, a multi-agent system may be required to perform tasks leveraging the information from the perception agents, for example, in a multi-UAV target tracking problem, a subset of UAVs may be tasked to ensure accurate tracking of the moving target, while other UAVs may be tasked to interdict the target given the target trajectory.
@@
Related work. This problem is closely related to the sensor placement problem, which aims to find an optimal subset of agents to participate in a perception task under limited sensing resources. Unlike traditional sensor coverage problems that emphasize spatial reach or visibility, the goal is to maximize the informativeness of the selected agents’ observations about the environment or latent variables of interest. Selecting too many agents wastes resources and increases redundancy, while selecting too few may lead to poor estimation or inference accuracy. To balance these trade-offs, many existing approaches formulate the objective using information-theoretic criteria, such as mutual information or entropy reduction Krause and Guestrin (2007); Singh et al. (2009). When the agents’ policies are finite, the environment is deterministic, and the inference objective is stationary—as in multi-robot path planning for predicting algae content in a lake—the resulting objectives often exhibit submodularity. This property enables the design of greedy selection algorithms with provable near-optimal performance guarantees and strong scalability for large multi-agent systems. However, in stochastic multi-agent systems, where the policy space is infinite (often parameterized by deep neural networks) and the unknown quantities to be inferred can be a stochastic process (such as tracking a moving agent), existing results do not directly apply.
@@
The problem of decentralized active sensing and perception has been studied extensively in multi-agent systems. Early work, such as Kreucher et al. (2005), formulated sensor management as a centralized active sensing problem, where a single decision maker plans sensing actions to maximize expected information gain using Bayesian filtering. While effective for small-scale systems, such centralized formulations are computationally expensive and scale poorly with the number of agents. Subsequent research has focused on decentralized and scalable planning methods. Satsangi et al. Satsangi et al. (2015, 2020) studied a class of partially observable Markov decision processes (POMDPs) whose value functions exhibit submodularity, and leveraged this property to design efficient algorithms for dynamic sensor selection with near-optimal guarantees. In contrast, Kumar et al. Kumar et al. (2017) considered multi-agent planning under stochastic dynamics, where submodularity appears in the reward structure rather than the value function. Their framework provides a theoretical foundation for decentralized decision making with submodular rewards, enabling agents to coordinate without centralized control. Best et al. Best et al. (2019) proposed Dec-MCTS, a decentralized Monte Carlo tree search method, for multi-robot active perception, which can handle stochasticity, but does not exploit the submodular structure of reward functions for efficiency or near-optimality guarantees. Lauri and Oliehoek Lauri and Oliehoek (2020) introduced a prediction-reward framework that quantifies the uncertainty in a centralized state estimate obtained after all agents complete sensing. To enable decentralized computation, they approximate this centralized reward using the expected accuracy of each agent’s prediction of the final belief. This reformulation expresses the global information objective as a standard Dec-POMDP reward function dependent only on local states, actions, and prediction actions, allowing the use of existing Dec-POMDP solvers while preserving the informativeness objective. However, this method also does not employ submodularity for more efficient computation.
@@
Recent research also studied the multi-robot multi-target tracking problem in deterministic, nonlinear systems Corah and Michael (2021). They employ the receding horizon planning framework and consider at each planning epoch, the team of robots aims to optimize the mutual information between the target states and the observations generated by a set of feasible trajectories for each robot. Because they consider a finite set of trajectories for each robot and deterministic dynamics, such an objective function can be shown to be submodular and monotone. By leveraging the property of a submodular function, they develop an algorithm to efficiently compute the near-optimal trajectories for multiple robots.
@@
Our contributions. Existing work on sensor selection or motion planning of active sensing agents considers finitely many sensors Singh et al. (2009) or discrete trajectories Corah and Michael (2021). In our setting, though the set of agents is finite, each agent is to compute an observation-based stochastic policy, and therefore, the policy space for agents is infinite. Therefore, existing GreedyMax algorithms for submodular Nemhauser et al. (1978) cannot be directly applied to compute jointly the subset of agents and their decentralized perception policies. Our problem can not be directly mapped to decentralized planning for multi-agent systems given submodular rewards because the information objectives cannot be decomposed into submodular reward given the joint state and joint actions (see Section 3). Existing Dec-POMDP for active perception Best et al. (2019); Lauri and Oliehoek (2020) does not solve the agent selection problem jointly with the policy synthesis.
@@
However, these results only enable us to use approximate algorithms for submodular maximization to solve the agent selection problem, assuming each selected agent follows a pre-defined decentralized policy. To achieve joint agent selection and policy synthesis, we propose the IMAS2 (Information-theoretic Multi-Agent Selection and Sensing) algorithm, a variant of the GreedyMax algorithm Nemhauser et al. (1978) to select one agent and determine its perception policy at a time based on the principle of maximizing the marginal gain. We then prove that, under additional constraints on subsequent maximal marginal gains, this algorithm can provide a strong (11/e)(1-1/e)-guarantee on the performance. The proposed algorithm achieves scalability and near-optimal performance, bridging rigorous submodular optimization theory with practical decentralized planning in multi-agent active perception. Finally, we formulate the agent selection and policy optimization problem as a two-layer optimization: an inner layer that computes an optimal perception policy, with respect to maximizing information gain of a chosen agent, in addition to observations of selected agents. And an outer layer efficiently selects $k$ agents to achieve the approximately optimal cooperative active perception. The results are then experimentally validated.
@@ 2. Preliminary
$f(S)+f(T)\geq f(S\cup T),$
holds for all $S,T\subset\Omega$. Define the first-order partial derivative (equivalently, the gain) of an element $j\notin S$ in the context $S$ as
$f(j|S)=f(S\cup\{j\})-f(S).$
The function $f$ is submodular Fujishige (2005) if for all $S,T\subset\Omega$, it holds that
$f(S)+f(T)\geq f(S\cup T)+f(S\cap T).$
@@
$f(j|S)\geq f(j|T)\quad\text{for all }S\subset T,\,j\notin T.$
Next, we introduce the definition of entropy and mutual information. Entropy is commonly employed to quantify the uncertainty about a random variable Khouzani and Malacaria (2017). The conditional entropy of a random variable $X_{2}$ given another random variable $X_{1}$ is defined by
@@ 3.1. Inferring Latent State Sequence
In this subsection, we study a class of active perception objectives for minimizing the uncertainty in the estimation of the joint state trajectories. In particular, we are interested in finding an (approximately)-optimal subset $\mathcal{K}$ of the agents to carry out the active perception tasks and computing their individual policies, for some $\mathcal{K}\subset\mathcal{N}$. The agents in $\mathcal{N}\setminus\mathcal{K}$ are excluded from the perception task and their observations do not contribute to the perception objective.
@@
See appendix. ∎
@@
Let $\Omega\coloneqq\{Y_{i},i\in\mathcal{N}\}$ be the set of observations for all agents. $Y_{A}\subset Y_{B}\subset\Omega$, and $Y_{j}\in\Omega\setminus Y_{B}$.
$I(X;Y_{B})-I(X;Y_{A})=I(X;Y_{B}\setminus Y_{A}|Y_{A})$
Since conditional mutual information is always non-negative Cover and Thomas (2006), $I(X;Y_{B})-I(X;Y_{A})\geq 0$ and thus the function $g(\cdot)$ is monotone. To show $g(\cdot)$ is submodular, we need to show that
$I(X;Y_{A}\cup\{Y_{j}\})-I(X;Y_{A})\geq I(X;Y_{B}\cup\{Y_{j}\})-I(X;Y_{B}).$
And based on the chain rule of mutual information,
$I(X;Y_{A}\cup\{Y_{j}\})-I(X;Y_{A})=I(X;Y_{j}|Y_{A}).$
$H(Y_{j}|Y_{A})\geq H(Y_{j}|Y_{B})$
$H(Y_{j}|Y_{A})-H(Y_{j}|X)\geq H(Y_{j}|Y_{B})-H(Y_{j}|X)$
$H(Y_{j}|Y_{A})-H(Y_{j}|X,Y_{A})\geq H(Y_{j}|Y_{B})-H(Y_{j}|X,Y_{B})$
$I(Y_{j};X|Y_{A})\geq I(Y_{j};X|Y_{B})$
$I(X;Y_{j}|Y_{A})\geq I(X;Y_{j}|Y_{B}),$
where the third step is due to Lemma 3 and the last step is because conditional mutual information is symmetric. ∎
@@
The above analysis is based on a pre-defined decentralized policy. It can be used for selecting a subset of agents whose observations are most informative for estimating the joint state trajectory. Leveraging the monotone, submodular property, the original GreedyMax algorithm Nemhauser et al. (1978) can find such a subset of agents with (11/e)(1-1/e) performance guarantee on the suboptimality.
@@ 3.3. Environment Secret Estimation
In this section, we consider a perception objective besides the aforementioned environment trajectory estimation. Consider a random variable $Z$ defined as a surjective function of $X_{e}$ i.e. , $Z=f(X_{e})$. It is noted that $H(X_{e})$ is a constant and hence $H(Z)$ is a constant.
@@
Lemma 0.
Under Assumption 2, let $A\subset\mathcal{N}$, $Y_{A}=\{Y_{i}:i\in A\}$, and let $Z=f(X_{e})$ denote the secret. Define
$g(Y_{A})\coloneqq I(Z;Y_{A}).$
Then the function $g(\cdot)$ is submodular and monotone.
@@
We first note that $h(Y_{A})\coloneqq I(X_{e};Y_{A}|Z)$ is submodular, by an argument analogous to the proof of Lemma 4, leveraging the property that $Y_{j}$ is conditionally independent from $Y_{i}$ given $X$ and thus $H(Y_{j}|X_{e},Y_{A},Z)=H(Y_{j}|X_{e})$.
Moreover, note that $I(X_{e};Y_{A}|Z)=H(X_{e}|Z)-H(X_{e}|Z,Y_{A}).$ Because $H(X_{e}|Z)$ is independent of $Y_{A}$ and $I(X_{e};Y_{A}|Z)$ is submodular, then $H(X_{e}|Z,Y_{A})$ is submodular. By equation (6),
$I(Z;Y_{A})=I(X_{e};Y_{A})-H(X_{e})+H(Z)+H(X_{e}|Z,Y_{A}),$
According to Lemma 8, $I(X_{e};Y_{A})$ is submodular. This result, combined with $H(X_{e}|Z,Y_{A})$ being submodular and $H(X_{e})$, $H(Z)$ being constants, we can derive that $I(Z;Y_{A})$ is submodular using the property that the sum of submodular functions is submodular.
Lastly, $I(Z;Y_{A})$ can be shown to be monotone in $Y_{A}$ by similar arguments in the proof of Lemma 4. ∎
@@ 4.1. Approximation Algorithm and Performance Guarantee
A key property of submodular functions is that they admit strong approximation guarantees when optimized under cardinality constraints. However, because we are simultaneously selecting the agent and determining their policy, the proof of approximate optimality needs to be modified. Next, we derive the performance guarantee of the IMAS2 algorithm. Let $({\mathcal{K}}^{\star},\bm{\pi}^{\star})$ be the optimal solution to (1). Let $({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})$ be the solution returned by Algorithm 1 after the $i$-th selection. By default, $\mathcal{K}^{(0)}=\varnothing$ and $\bm{\pi}^{(0)}=\varnothing$. Let $f(A,\bm{\pi}_{A})=I(X;Y_{A},\mathcal{M}_{\bm{\pi}_{A}})$ for any subset $A\subset\mathcal{A}$ and $\bm{\pi}_{A}=\{\pi_{i},i\in A\}$ (or $f(A,\bm{\pi}_{A})=I(X_{e};Y_{A},\mathcal{M}_{\bm{\pi}_{A}})$ for (2), and $f(A,\bm{\pi}_{A})=I(Z;Y_{A},\mathcal{M}_{\bm{\pi}_{A}})$ for (5)). By default, $f(\varnothing,\varnothing)=0$.
@@
Let $\Delta_{i}\coloneqq f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})-f({\mathcal{K}}^{(i-1)},\bm{\pi}^{(i-1)})$ for $1\leq i<k$. It holds that
$\Delta_{i+1}\leq\Delta_{i},\forall 0\leq i<k.$
@@
Recall $(j^{\ast},\pi_{j^{\ast}})$ is the choice of agent and its policy selected at the $i$-th iteration of Algorithm 1. According to Algorithm 1,
$-f({\mathcal{K}}^{(i-2)},\bm{\pi}^{(i-2)})\big)$
$=$ $f({\mathcal{K}}^{(i-1)},\bm{\pi}^{(i-1)})-f({\mathcal{K}}^{(i-2)},\bm{\pi}^{(i-2)})$
$=$ $\Delta_{i}.$
Here, $(i)$ follows from the submodularity of the objective function: the marginal contribution of agent $j^{\ast}$ using policy $\pi_{j^{\ast}}$ with respect to $(\mathcal{K}^{(i-2)},\bm{\pi}^{(i-2)})$ is greater than or equal to its contribution with respect to $(\mathcal{K}^{(i-1)},\bm{\pi}^{(i-1)})$ under the same policy. And $(ii)$ is due to the greedy choice made at step i1i-1. ∎
@@
For each $i\geq 1$, let the gap between the value under the optimal solution and the value under the solution returned by the $i$-th selection be $g_{i}\coloneqq f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})$.
$\overset{(ii)}{=}$ $k\Delta_{i+1}-\Delta_{i+2}$
$=$ $\;k\Delta_{i+2}-(k+1)\Delta_{i+2}+k\Delta_{i+1}$
$\overset{(iii)}{\leq}$ $k\Delta_{i+2}.$
Using successive gaps, $\Delta_{i+1}\geq\frac{1}{k}g_{i}$ and $g_{i+1}=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i+1)},\bm{\pi}^{(i+1)})=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})-f({\mathcal{K}}^{(i+1)},\bm{\pi}^{(i+1)})+f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})-f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})=g_{i}-\Delta_{i+1}\leq g_{i}-\frac{1}{k}g_{i}=(1-\frac{1}{k})g_{i}$. (This part of the proof is similar to the original proof in Nemhauser et al. (1978)).
Consequentially, $g_{k}\leq(1-\frac{1}{k})^{k}g_{0}$ where $g_{0}=f({\mathcal{K}}^{\star},\bm{\pi}^{\star})$. Using the exponential bound, $(1-\frac{1}{k})^{k}\leq e^{-1}$ for any $k\geq 1$. Equation (7) is then derived. ∎

### v2

@@ Abstract.
Decentralized Partially Observable Markov Decision Process (Dec-POMDP); Agent Selection; Active Perception; Submodular Optimization.
doi:
STZK9664
@@ 1. Introduction
Autonomous multi-agent systems are increasingly deployed in environments where agents must actively gather information under uncertainty. Applications include teams of robots for surveillance and search-and-rescue Rybski2000surveillance; Kantor2006; Rosenfeld2015Intelligent, sensor networks for target tracking Tian2006Target; Jing2020Pose, and cooperative perception in autonomous driving and aerial monitoring Xu2020hitmac; wang2023corecooperativereconstructionmultiagent. In such settings, an active perception agent must decide not only how to act, but also what to sense, given the past observation. In multi-agent cooperative active perception, it must additionally account for its knowledge of other agents’ possible observations to avoid redundancy and achieve better coordination. Meanwhile, in addition to perception, a multi-agent system may be required to perform tasks leveraging the information from the perception agents, for example, in a multi-UAV target tracking problem, a subset of UAVs may be tasked to ensure accurate tracking of the moving target, while other UAVs may be tasked to interdict the target given the target trajectory.
@@
Related work. This problem is closely related to the sensor placement problem, which aims to find an optimal subset of agents to participate in a perception task under limited sensing resources. Unlike traditional sensor coverage problems that emphasize spatial reach or visibility, the goal is to maximize the informativeness of the selected agents’ observations about the environment or latent variables of interest. Selecting too many agents wastes resources and increases redundancy, while selecting too few may lead to poor estimation or inference accuracy. To balance these trade-offs, many existing approaches formulate the objective using information-theoretic criteria, such as mutual information or entropy reduction Krause2007Near; Singh2009informative. When the agents’ policies are finite, the environment is deterministic, and the inference objective is stationary—as in multi-robot path planning for predicting algae content in a lake—the resulting objectives often exhibit submodularity. This property enables the design of greedy selection algorithms with provable near-optimal performance guarantees and strong scalability for large multi-agent systems. However, in stochastic multi-agent systems, where the policy space is infinite (often parameterized by deep neural networks) and the unknown quantities to be inferred can be a stochastic process (such as tracking a moving agent), existing results do not directly apply.
@@
The problem of decentralized active sensing and perception has been studied extensively in multi-agent systems. Early work, such as Kreucher2005Sensor, formulated sensor management as a centralized active sensing problem, where a single decision maker plans sensing actions to maximize expected information gain using Bayesian filtering. While effective for small-scale systems, such centralized formulations are computationally expensive and scale poorly with the number of agents. Subsequent research has focused on decentralized and scalable planning methods. Satsangi et al. Satsangi_Whiteson_Oliehoek_2015; Satsangi2020Active studied a class of partially observable Markov decision processes (POMDPs) whose value functions exhibit submodularity, and leveraged this property to design efficient algorithms for dynamic sensor selection with near-optimal guarantees. In contrast, Kumar et al. Kumar2017Decentralized considered multi-agent planning under stochastic dynamics, where submodularity appears in the reward structure rather than the value function. Their framework provides a theoretical foundation for decentralized decision making with submodular rewards, enabling agents to coordinate without centralized control. Best et al. Best2019MCTS proposed Dec-MCTS, a decentralized Monte Carlo tree search method, for multi-robot active perception, which can handle stochasticity, but does not exploit the submodular structure of reward functions for efficiency or near-optimality guarantees. Lauri and Oliehoek Lauri2020prediction introduced a prediction-reward framework that quantifies the uncertainty in a centralized state estimate obtained after all agents complete sensing. To enable decentralized computation, they approximate this centralized reward using the expected accuracy of each agent’s prediction of the final belief. This reformulation expresses the global information objective as a standard Dec-POMDP reward function dependent only on local states, actions, and prediction actions, allowing the use of existing Dec-POMDP solvers while preserving the informativeness objective. However, this method also does not employ submodularity for more efficient computation.
@@
Recent research also studied the multi-robot multi-target tracking problem in deterministic, nonlinear systems corahScalableDistributedPlanning2021. They employ the receding horizon planning framework and consider at each planning epoch, the team of robots aims to optimize the mutual information between the target states and the observations generated by a set of feasible trajectories for each robot. Because they consider a finite set of trajectories for each robot and deterministic dynamics, such an objective function can be shown to be submodular and monotone. By leveraging the property of a submodular function, they develop an algorithm to efficiently compute the near-optimal trajectories for multiple robots.
@@
Our contributions. Existing work on sensor selection or motion planning of active sensing agents considers finitely many sensors Singh2009informative or discrete trajectories corahScalableDistributedPlanning2021. In our setting, though the set of agents is finite, each agent is to compute an observation-based stochastic policy, and therefore, the policy space for agents is infinite. Therefore, existing GreedyMax algorithms for submodular nemhauser1978analysis cannot be directly applied to compute jointly the subset of agents and their decentralized perception policies. Our problem can not be directly mapped to decentralized planning for multi-agent systems given submodular rewards because the information objectives cannot be decomposed into submodular reward given the joint state and joint actions (see Section 3). Existing Dec-POMDP for active perception Best2019MCTS; Lauri2020prediction does not solve the agent selection problem jointly with the policy synthesis.
@@
However, these results only enable us to use approximate algorithms for submodular maximization to solve the agent selection problem, assuming each selected agent follows a pre-defined decentralized policy. To achieve joint agent selection and policy synthesis, we propose the IMAS2 (Information-theoretic Multi-Agent Selection and Sensing) algorithm, a variant of the GreedyMax algorithm nemhauser1978analysis to select one agent and determine its perception policy at a time based on the principle of maximizing the marginal gain. We then prove that, under additional constraints on subsequent maximal marginal gains, this algorithm can provide a strong (11/e)(1-1/e)-guarantee on the performance. The proposed algorithm achieves scalability and near-optimal performance, bridging rigorous submodular optimization theory with practical decentralized planning in multi-agent active perception. Finally, we formulate the agent selection and policy optimization problem as a two-layer optimization: an inner layer that computes an optimal perception policy, with respect to maximizing information gain of a chosen agent, in addition to observations of selected agents. And an outer layer efficiently selects $k$ agents to achieve the approximately optimal cooperative active perception. The results are then experimentally validated.
Assumptions and Scope. We focus on Dec-POMDPs with observation independent sensing models, formalized by Assumption 1. Importantly, Subsection 3.1 (Inferring Latent State Sequence) only requires this observation conditional independence and does not require transition independence. In contrast, Subsection 3.2 (Inferring Environment State Sequence) and Section 3.3 (Environment Secret Estimation) adopt the stronger transition- and observation independence conditions in Assumption 2. We highlight these assumptions explicitly to distinguish the settings covered by each result.
@@ 2. Preliminary
$f(U)+f(V)\geq f(U\cup V),$
holds for all $U,V\subset\Omega$. Define the first-order partial derivative (equivalently, the gain) of an element $j\notin U$ in the context $U$ as
$f(j\mid U)=f(U\cup\{j\})-f(U).$
The function $f$ is submodular fujishige2005submodular if for all $U,V\subset\Omega$, it holds that
$f(U)+f(V)\geq f(U\cup V)+f(U\cap V).$
@@
$f(j\mid U)\geq f(j\mid V)\quad\text{for all }U\subset V,\,j\notin V.$
Next, we introduce the definition of entropy and mutual information. Entropy is commonly employed to quantify the uncertainty about a random variable khouzani2017leakage. The conditional entropy of a random variable $X_{2}$ given another random variable $X_{1}$ is defined by
@@ 3.1. Inferring Latent State Sequence
In this subsection, we study a class of active perception objectives for minimizing the uncertainty in the estimation of the joint state trajectories. In particular, we are interested in finding an (approximately)-optimal subset $\mathcal{K}$ of the agents to carry out the active perception tasks and computing their individual policies, for some $\mathcal{K}\subset\mathcal{N}$. The agents in $\mathcal{N}\setminus\mathcal{K}$ are excluded from the perception task and their observations do not contribute to the perception objective. Importantly, selecting a subset $\mathcal{K}$ does not remove the other agents from the system dynamics: all agents in $\mathcal{N}$ still execute their respective policies and evolve under the original joint transition model. The only change is that the inference objective is evaluated using the collective observations of agents in $\mathcal{K}$ (e.g., for $\mathcal{N}=\{1,\ldots,5\}$ and $\mathcal{K}=\{1,2,4\}$, all five agents act and affect the state evolution, while only $Y_{1},Y_{2},Y_{4}$ are used for estimating $S_{0:T}$).
@@
See appendix [^1] You can find the appendix in https://github.com/AronYoung414/multi-agent-active-perception-grid-world.. ∎
@@
See appendix. ∎
@@
The above analysis is based on a pre-defined decentralized policy. It can be used for selecting a subset of agents whose observations are most informative for estimating the joint state trajectory. Leveraging the monotone, submodular property, the original GreedyMax algorithm nemhauser1978analysis can find such a subset of agents with (11/e)(1-1/e) performance guarantee on the suboptimality.
@@ 3.2. Inferring Environment State Sequence
Moreover, under the transition- and observation-independence conditions below, the Dec-POMDP induced by selecting a subset of agents can be constructed in a standard way by restricting the dynamics and observation model to the environment state and the selected agents’ local states. In particular, because the joint transition and observation models factor across agents, the evolution of the selected agents depends only on their own local states/actions and the environment process, and the non-selected agents can be marginalized without affecting the resulting reduced model.
@@ 3.3. Environment Secret Estimation
In this section, we consider a perception objective besides the aforementioned environment trajectory estimation. As in Subsection 3.2, our analysis for this setting assumes the transition- and observation-independence conditions in Assumption 2, so that the Dec-POMDP induced by a reduced agent set can be constructed by restricting to the environment state and the selected agents’ local states/observations. Consider a random variable $Z$ defined as a surjective function of $X_{e}$ i.e. , $Z=f(X_{e})$. It is noted that $H(X_{e})$ is a constant and hence $H(Z)$ is a constant.
@@
Lemma 0 (Approximate Submodularity).
Let
$g(A)\coloneqq I(Z;Y_{A}),$
where $Z$ is a secret variable and $Y_{A}$ denotes the joint observations collected by agent set $A$. Then $g(\cdot)$ is monotone and $\epsilon$-approximately submodular, i.e., there exists a submodular function
$h(A)\coloneqq I(X_{e};Y_{A})$
such that for all $A\subseteq\mathcal{N}$,
$(1-\epsilon)h(A)\leq g(A)\leq(1+\epsilon)h(A),$
where
$\epsilon\coloneqq\max_{A}\frac{I(X_{e};Y_{A}\mid Z)}{I(X_{e};Y_{A})}\in(0,1].$
@@
See appendix. ∎
@@ 4.1. Approximation Algorithm and Performance Guarantee
A key property of submodular functions is that they admit strong approximation guarantees when optimized under cardinality constraints. However, because we are simultaneously selecting the agent and determining their policy, the proof of approximate optimality needs to be modified. Next, we derive the performance guarantee of the IMAS2 algorithm when the objective functions are submodular.
In (1), let $f(A,\bm{\pi}_{A})=I(X;Y_{A},\mathcal{M}_{\bm{\pi}_{A}})$ and in (2), let $f(A,\bm{\pi}_{A})=I(X_{e};Y_{A},\mathcal{M}_{\bm{\pi}_{A}})$, for any subset $A\subset\mathcal{A}$ and $\bm{\pi}_{A}=\{\pi_{i},i\in A\}$. Let $({\mathcal{K}}^{\star},\bm{\pi}^{\star})$ be the optimal solution to (1) (or (2)). Let $({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})$ be the solution returned by Algorithm 1 after the $i$-th selection. By default, $\mathcal{K}^{(0)}=\varnothing$ and $\bm{\pi}^{(0)}=\varnothing$. By default, $f(\varnothing,\varnothing)=0$.
@@
Let $\Delta_{i}\coloneqq f({\mathcal{K}}^{(i)},\bm{\pi}^{(i)})-f({\mathcal{K}}^{(i-1)},\bm{\pi}^{(i-1)})$ for $1\leq i<k$. It holds that $\Delta_{i+1}\leq\Delta_{i},\forall 0\leq i<k.$
@@
See appendix. ∎
@@
See appendix. ∎
Remark 0.
Since the objective function in (5) is only approximately submodular, analyzing the performance of the proposed algorithm requires explicitly accounting for the approximation ratio $\epsilon$. We defer a performance analysis and the establishment of formal guarantees for the secret inference case to future work.